OpenStudy (anonymous):

7/3 divided by (3/5 + 1/10) = order of operations

OpenStudy (ddcamp):

Start with parentheses: 3/5 + 1/10 = 6/10 + 1/10 = 7/10 Then, to divide 7/3 by 7/10, multiply by the reciprocal: 7/3 ÷ 7/10 = 7/3 × 10/7 = 70/21 = 10/3

OpenStudy (anonymous):

it says type a mixed number?

OpenStudy (ddcamp):

10/3 = 9/3 + 1/3 = 3 + 1/3 = 3 1/3

OpenStudy (anonymous):

thanks can you help me with more

OpenStudy (ddcamp):

Sure

OpenStudy (anonymous):

order of operations again. (1/3 + 1/15) x 1/3=

OpenStudy (anonymous):

simplify your answer.

OpenStudy (anonymous):

i got this one wrong

OpenStudy (ddcamp):

Again, start in the parentheses by finding the lowest common denominator (15): 5/15 + 1/15 → 6/15 (6/15)×1/3 = 6 / 30 = 1/5

hero (hero):

Well, actually, you have to do order of operations, then convert the result to a mixed fraction.

hero (hero):

\[\frac{7}{3} \div \left(\frac{3}{5} + \frac{1}{10}\right) \\ =\frac{7}{3} \div \left(\frac{6}{10} + \frac{1}{10}\right) \\=\frac{7}{3} \div \left(\frac{7}{10} \right) \\=\frac{7}{3} \times \frac{10}{7} \\=\frac{10}{3} \\ = \frac{9 + 1}{3} \\= \frac{9}{3} + \frac{1}{3} \\= 3 + \frac{1}{3}\]